OpenStudy (anonymous):

I need help on this question :p Let f=(x)= 3x + 2 and g(x) = 6x – 7. Find f=(x)– g(x).

OpenStudy (binary3i):

f(x)= 3x+2 g(x)=6x-7 f(x)-g(x)= 3x+2-(6x-7)=3x-6x+2+7=-3x+9 so the answer is b

OpenStudy (anonymous):

Thanks :)